# ChronoRoot 2.0: An Open AI-Powered Platform for 2D Temporal Plant Phenotyping

### v1

@@ 3.1.1 Multi-Class Segmentation Performance
Building upon these improvements in binary segmentation, we evaluated the nnUNet’s performance in discriminating among six distinct plant structures. This multi-class capability represents a significant advancement over the original system, enabling tracking of multiple plant organs throughout development.
The boxplots in Figure 3 illustrate the distribution of segmentation metrics across all evaluated structures. Dice coefficients (Figure 3A) show consistent performance across different plant organs, with expected variation in median values reflecting the relative difficulty in segmenting each structure. Hausdorff distances (Figure 3B) provide insight into the spatial accuracy of the segmentation boundaries, with lower values indicating more precise delineation of plant structures. Most notably, despite their complex morphology, root structures maintained high segmentation accuracy while preserving fine structural details necessary for subsequent analysis.
Beyond raw segmentation accuracy, successful root system analysis depends critically on preserving key morphological traits. We therefore conducted additional validation of the skeletonized root segmentations using completeness and correctness metrics, which directly assess structural fidelity [16].
Completeness measures the extent to which the extracted skeleton retains the original root structure, with higher values indicating fewer missing segments. This ensures that the segmentation does not omit fine root details, which are essential for accurate architectural quantification. The main and lateral roots exhibited high completeness (0.951 ± 0.060 and 0.919 ± 0.112, respectively), suggesting minimal loss of structural information.
Correctness evaluates the presence of extraneous or spurious branches in the extracted skeleton. A high correctness score indicates that the segmentation accurately follows the true root architecture without introducing artifacts or unintended branches. The correctness values for main and lateral roots (0.963 ± 0.050 and 0.935 ± 0.119, respectively) confirm that the extracted skeletons maintain structural fidelity and avoid unnecessary deviations.
Figure 3: Multi-class segmentation performance of ChronoRoot 2.0. Segmentation quality measured by Dice coefficient (A, higher is better) and Hausdorff distance (B, lower is better, in mm) across plant structures. Root-specific metrics (C, completeness and correctness, higher is better) demonstrate structural fidelity preservation.
@@ 3.2 Demonstration of Software Capabilities Through Use Cases
To validate ChronoRoot 2.0’s practical utility across diverse experimental scenarios, we implemented three use cases. The first one examines root system architecture under long day and continuous light condition following the original publication [6], while the second analyzes published data from the transcription factor gene NF-YA10 over-expressing plants [17]. The final use case demonstrates the high-throughput screening capabilities of the system, on an etiolation experiment. Importantly, all figure subpanels presented in these use cases are direct outputs from ChronoRoot 2.0, with the only modification being the addition of asterisks (*) to indicate statistical differences. The statistical analysis values are also provided by the software as accompanying text files, requiring no additional analysis beyond what the software automatically generates.
@@
Figure 4: Use Case 1 - Root system architecture dynamics under different light conditions. Comparison of long-day (16h/8h, blue) versus continuous light (24h, orange) shows divergent growth patterns. A - All basic Archictural RSA parameters, B - FPCA analysis of Main Root Length, C - Fourier transform of Main Root Growth Rate, D - FPCA analysis of Lateral Root Length. Error bands: standard error.
@@
Plant materials: All plants used in this study are in Columbia-0 background. pNF-YA10:GFP-NF-YA10miRres (NF-YA10miRres) stable lines were obtained by transforming Arabidopsis plants with a construct bearing 2000-pb region upstream of the start codon of NF-YA10 amplified from genomic DNA (promoter region) and the coding sequence (CDS) of NF-YA10 without miRNA cleavage site amplified from cDNA, thus resisting miR169-mediated post-transcriptional silencing of NF-YA10 mRNA. More details were published at [17].
@@
Figure 5: Use Case 2 - Altered root architecture in NF-YA10miRres plants. A - Qualitative analysis of convex hull showing root system coverage differences between genotypes. B - Quantitative analysis of convex hull metrics (area and aspect ratio) between NF-YA10miRres plants (red) and Col0 controls (blue). C - Quantitative analysis of average emergence angle and base-tip angle for the first lateral root, demonstrating consistently wider angles in NF-YA10miRres plants compared to Col0. Error bars: standard error.
@@ 3.2.3 Use Case 3 - High-throughput Analysis of Etiolation in Arabidopsis thaliana seedlings
Figure 6: Use Case 3 - High-throughput analysis of seedling etiolation. A - Representative infrared images showing temporal progression of etiolated seedling development. B - Hypocotyl, root and total area measurements with their corresponding growth rates. C - Germination curves showed T50 at approximately 13 hs after light stimulation with no significant differences between the analyzed genotypes. D - Comparison of Manual (blue) and automatic (orange) hypocotyl length determinations at 36, 48 and 60 hours after light stimulation (no significant difference observed). Error bars: standard error.
@@ 4 Discussion
The multi-class segmentation approach addresses a significant challenge in developmental studies by enabling simultaneous analysis of multiple plant structures. Although root growth on agar plates differs from soil conditions, the ability to precisely track both below and above-ground organs provides valuable insights into developmental coordination. The etiolation response study demonstrates how this capability can reveal resource allocation patterns during early development, with the simultaneous tracking of hypocotyl elongation and root growth providing a more complete understanding of seedling responses to dark conditions. Importantly, while the original ChronoRoot 2.0 multi-class segmentation model has been designed to identify plant organs for Arabidopsis thaliana, the use of nnUNet as segmentation model makes it easy its re-training for other plant species.
@@
Several technical considerations should guide the application of these tools. Image quality and plant positioning remain critical factors for optimal segmentation performance, particularly in high-throughput scenarios. While the system can process up to 100 seedlings per plate, careful experimental design is needed to minimize plant overlap during growth. Additionally, the standardized growth conditions required for imaging consistency should be considered when interpreting biological findings.
The analysis frameworks implemented in ChronoRoot 2.0 open new possibilities for understanding plant development, even within the constraints of traditional growth systems. The ability to automatically quantify subtle architectural differences and temporal patterns enables systematic comparison of developmental responses across genotypes and conditions. These capabilities are particularly valuable for studies investigating the genetic and environmental regulation of plant development, where precise quantification of phenotypic differences is essential. Furthermore, high-throughput comprehensive phenotyping emerges as a powerful tool for genome-wide association studies (GWAS) and the identification of key genes participating in plant development.
@@
Acknowledgments
Author Contributions
@@ Acknowledgments
  • F. Ariel, A. Barrios, F. Catulo, and T. Blein conducted the biological experiments analyzed in the study
  • F. S. Rodriguez, F. E. Aballay, F. B. Catulo, M. F. Mammarella, M. F. Legascue, and S. N. Villarreal performed manual image annotation for training the nnUNet module
@@
Funding
Conflicts of Interest
The authors declare that there is no conflict of interest regarding the publication of this article.
Data Availability
  • The annotated image dataset used for training and validation contains 693 infrared images of Arabidopsis thaliana seedlings with expert annotations for six distinct structural classes. This dataset, along with the trained nnUNet weights, is publicly available without restrictions at https://huggingface.co/datasets/ngaggion/ChronoRoot_nnUNet. The dataset includes both raw images and their corresponding multi-class segmentation masks in .png format, following nnUNet’s standardized structure for 2D images.
  • The complete source code of ChronoRoot 2.0, including the implementation of all analysis methods described in this paper, is freely available under the GNU General Public License v3.0 at https://github.com/ngaggion/ChronoRoot2. This repository contains the full software implementation and comprehensive documentation to set up the system and utilizing the software.
Supplementary Materials
@@
Figure S4: Use Case 3 - FPCA Analysis.To complement the growth dynamics presented in Figure 6-B, we performed FPCA to summarize and compare the developmental trajectories of hypocotyl length, total root length, and projected area across genotypes G1, G2, and G3. The first two principal components accounted for over 90% of the total variance in each trait, capturing the main temporal patterns of growth. For each metric, the top panels show the mean trajectories with standard error bars for each genotype. The middle and bottom rows illustrate the distribution of FPCA scores by group for PC1 and PC2, along with visual interpretations of each component.

### v2

@@ 3.1.1 Multi-Class Segmentation Performance
Building upon these improvements in binary segmentation, we evaluated the nnUNet’s performance in discriminating among six distinct plant structures. This multi-class capability represents a significant advancement over the original system, enabling tracking of multiple plant organs throughout development. The dataset was partitioned, within each of the three major experimental categories (etiolation, germination, and plant root analysis), into training (70%), validation (10%), and test (20%) sets following a video-based splitting strategy to prevent data leakage.
Beyond standard segmentation overlap metrics, successful root system analysis depends critically on preserving key morphological traits. We therefore evaluated the skeletonized root segmentations using completeness and correctness metrics, which directly assess structural fidelity [17]. Completeness measures the extent to which the extracted skeleton retains the original root structure, with higher values indicating fewer missing segments. Correctness evaluates the presence of extraneous or spurious branches in the extracted skeleton, with high values indicating that the segmentation accurately follows the true root architecture without introducing artifacts.
Table 3 presents the overall performance averaged across all plant structures, and the completeness and correctness calculated for the complete root, for each model configuration. All variants achieved similar Dice coefficients, with the standard and residual architectures showing no significant differences between them, but both significantly outperforming their respective fast (non-TTA) variants according to Wilcoxon Pair Ranked Test. Processing times ranged from  1 to  5 seconds per image, with the fast variants providing approximately 3-fold speedup. Detailed per plant organ values are shown in Supplementary Table S1.
3.1.2 Across species generalization: Tomato.
To evaluate the generalizability of our approach to other plant species, we trained and tested nnUNet models on a tomato dataset. The data was partitioned by experimental setup, with one complete acquisition (24 plates) reserved for testing, resulting in 299 training images and 181 test images. Note that in the tomato dataset, leaves and petioles were annotated as a single combined aerial part class, reflecting species-specific morphological differences from Arabidopsis thaliana.
We evaluated two training strategies: (1) models trained exclusively on tomato data, and (2) models trained on combined tomato and Arabidopsis thaliana datasets. For the combined training approach, Arabidopsis annotations were preprocessed to match the tomato class structure by merging leaf and petiole classes into a single aerial part category. All four model configurations (Standard, Standard Fast, Residual, and Residual Fast) were evaluated under both training regimes.
The results (Table 4) reveal several important findings. First, residual architectures consistently outperformed standard architectures across all metrics, with particularly notable improvements in Hausdorff distance and correctness measures. Second, incorporating Arabidopsis training data significantly enhanced performance across all model configurations, with the combined training strategy yielding the best results. These findings demonstrate both the transferability of knowledge across plant species and the value of diverse training data for robust segmentation performance. Detailed per plant organ values are shown in Supplementary Table S2.
Table 4: Cross-species generalization and training strategy evaluation for the tomato dataset. Residual architectures show better results in Hausdorff Distance (HD) and Correctness (Cr), in the separated test set (n=181). The multi-species training strategy (Both) consistently outperforms training only on tomato data, showing that data diversity improves results across different morphologies.
Training Configuration Dice HD Cp Cr Time (s)
Tomato Standard 0.815±0.218 19.793±20.590 0.920±0.133 0.779±0.230 2.198
Tomato Standard (Fast) 0.801±0.220 24.357±22.973 0.910±0.157 0.733±0.253 0.813
Tomato Residual 0.843±0.200 15.430±18.900 0.908±0.153 0.868±0.189 5.290
Tomato Residual (Fast) 0.829±0.206 17.295±19.492 0.904±0.159 0.853±0.197 1.586
Both Standard 0.828±0.212 19.803±20.335 0.914±0.141 0.816±0.223 2.199
Both Standard (Fast) 0.822±0.214 19.722±19.897 0.908±0.156 0.791±0.232 0.768
Both Residual 0.863±0.201 11.089±16.640 0.916±0.138 0.896±0.177 5.310
Both Residual (Fast) 0.858±0.195 11.553±16.174 0.905±0.165 0.899±0.172 1.557
@@ 3.2 Demonstration of Software Capabilities Through Use Cases
To validate ChronoRoot 2.0’s practical utility across diverse experimental scenarios, we implemented three use cases. The first one examines root system architecture under long day and continuous light condition following the original publication [6], while the second analyzes published data from the transcription factor gene NF-YA10 over-expressing plants [18]. The final use case demonstrates the high-throughput screening capabilities of the system, on an etiolation experiment. Importantly, all figure subpanels presented in these use cases are direct outputs from ChronoRoot 2.0 and serve as representative examples of the automated reports generated when users analyze their own data or the provided demo datasets. The only modification to these outputs is the addition of asterisks to indicate statistical significance: * for p<0.05 and ** for p<0.001. While the figures prioritize visual clarity for phenotype comparison, the exact statistical analysis values—including p-values, means, and standard deviations—are provided by the software as accompanying text files within the output folders, requiring no additional analysis beyond what the software automatically generates.
Figure 3: Use Case 1 - Arabidopsis thaliana Root system architecture dynamics under different light conditions. Comparison of long-day (16h/8h, blue, n=23) versus continuous light (24h, orange, n=21) shows divergent growth patterns. A - All basic Architectural RSA parameters, B - FPCA analysis of Main Root Length, significative differences found in PC2 (p-value<0.001), C - Fourier transform of Total Root Growth Rate (significative differences found at both 24h and 12h periods, p-value<0.05), D - FPCA analysis of Lateral Root Length, significative differences found in PC2 (p-value<0.05). Error bands: standard error.
@@
Plant materials: All plants used in this study are in Columbia-0 background. pNF-YA10:GFP-NF-YA10miRres (NF-YA10miRres) stable lines were obtained by transforming Arabidopsis plants with a construct bearing 2000 bp region upstream of the start codon of NF-YA10 amplified from genomic DNA (promoter region) and the coding sequence (CDS) of NF-YA10 without miRNA cleavage site amplified from cDNA, thus resisting miR169-mediated post-transcriptional silencing of NF-YA10 mRNA. More details were published at [18].
@@
Figure 4: Use Case 2 - Altered root architecture in Arabidopsis thaliana NF-YA10miRres plants. A - Qualitative analysis of convex hull showing root system coverage differences between genotypes. B - Quantitative analysis of convex hull metrics (area and aspect ratio) between NF-YA10miRres plants (1 in orange, n=26), 2 in green, n=28) and Col0 controls (blue, n=16). C - Quantitative analysis of average emergence angle and base-tip angle for the first lateral root, demonstrating consistently wider angles in NF-YA10miRres plants compared to Col0. Error bars: standard error.
@@ 3.2.3 Use Case 3 - High-throughput Analysis of Etiolation in Arabidopsis thaliana seedlings
Figure 5: Use Case 3 - High-throughput analysis of Arabidopsis thaliana seedling etiolation. A - Representative infrared images showing temporal progression of etiolated seedling development. B - Hypocotyl, root and total area measurements with their corresponding growth rates (G1, blue, n=189. G2, orange, n=211; G3, green, n=180). C - Germination curves showed T50 at approximately 13 hs after light stimulation with no significant differences between the analyzed genotypes. D - Comparison of Manual (blue) and automatic (orange) hypocotyl length determinations at 36, 48 and 60 hours after light stimulation (no significant difference observed). Error bars: standard error.
@@
3.2.4 Use Case 4 - Multi-species Capability: Tomato Analysis
Plant materials: Tomato seeds of cultivar M82 were surface sterilized and sown on MS medium under standard conditions. To accommodate the larger size of tomato seedlings compared to Arabidopsis, plant density was reduced to two seeds per plate.
To further demonstrate the species-agnostic design of ChronoRoot 2.0, we applied the complete analysis pipeline to tomato seedlings, which present larger organs, thicker roots, and increased curvature compared to Arabidopsis. The nnUNet-based segmentation accurately identified and tracked main roots, lateral roots, and hypocotyls over time in both wild type and mutant plants (Figure 6-A), enabling the extraction of standard architectural and temporal traits without parameter tuning.
Quantitative analysis revealed marked differences between the two genotypes. Spatial descriptors such as accumulated root traces and convex hulls showed that mutant seedlings explored a substantially smaller area than wild type plants (Figure 6-B). Temporal measurements of main root, lateral root, hypocotyl, and total length highlighted an early and persistent reduction in growth in the mutant condition (Figure 6-D). Principal component analysis of the extracted traits captured most of the variance with the first two components and clearly separated wild type and mutant populations, reflecting differences in overall growth magnitude and temporal progression (Figure 6-C).
Figure 6: Use Case 4 - ChronoRoot 2.0 multi-species analysis on tomato seedlings. A - Representative time-lapse images of wild type (WT, left, n=5) and mutant (right, n=5) tomato seedlings, with nnUNet-based segmentation and organ tracking overlays. B - Spatial characterization of root system architecture showing accumulated root traces, last-day convex hull contours, and occupied growth area for mutant (top) and WT (bottom). C - Multivariate analysis of extracted traits: main root length dynamics, PCA variance explanation, boxplots for PC1 and PC2, and temporal interpretation of principal components. D - Organ-specific temporal growth curves for main root, lateral roots, hypocotyl, and total length, comparing WT and mutant seedlings.
@@ 4 Discussion
The multi-class segmentation approach addresses a significant challenge in developmental studies by enabling simultaneous analysis of multiple plant structures. Although root growth on agar plates differs from soil conditions, the ability to precisely track both below and above-ground organs provides valuable insights into developmental coordination. The etiolation response study demonstrates how this capability can reveal resource allocation patterns during early development, with the simultaneous tracking of hypocotyl elongation and root growth providing a more complete understanding of seedling responses to dark conditions. To demonstrate the system’s adaptability beyond Arabidopsis thaliana, we incorporated a tomato (Solanum lycopersicum) dataset. This addition validates that the self-configuring nnUNet core can be effectively retrained to handle the more robust and diverse morphologies of crop species. While applying the system to plants with fundamentally different architectures would require new annotated training data, this framework ensures that the adaptation process remains accessible to researchers without deep machine learning expertise.
While the software architecture is fundamentally modality-agnostic and offers a modular pathway for adaptation to other 2D imaging platforms (such as SPIRO [15]), the platform’s robustness is anchored in its integration with our custom hardware. By coupling the software with an open-source, affordable, and easily assembled hardware unit, we ensure high data quality and stable temporal resolution (15-minute intervals) without presenting a significant financial barrier to adoption.
To maximize the utility of these segmentation capabilities within the spatial limits of 12x12 cm plates, we implemented a dual-interface system designed to handle specific experimental constraints. The Standard Interface is designed for high-precision architectural tracking of normally 4-6 plants per plate. These experiments are typically limited to 10-14 days for Arabidopsis or 5-7 days for tomato, concluding when the main root reaches the plate bottom or grows along the surface invisible to imaging. Conversely, the Screening Interface accommodates up to 100 plants per plate but limits analysis to the early developmental window (typically 3-5 days). This mode allows for massive data collection before plant crowding and physical contact prevent reliable segmentation.
@@
The analysis frameworks implemented in ChronoRoot 2.0 open new possibilities for understanding plant development, even within the constraints of traditional growth systems. The ability to automatically quantify subtle architectural differences and temporal patterns enables systematic comparison of developmental responses across genotypes and conditions. These capabilities are particularly valuable for studies investigating the genetic and environmental regulation of plant development, where precise quantification of phenotypic differences is essential. Furthermore, high-throughput comprehensive phenotyping emerges as a powerful tool for genome-wide association studies and the identification of key genes participating in plant development.
@@
5 Code and Data Availability
  • The complete source code of ChronoRoot 2.0, including the implementation of all analysis methods described in this paper, is freely available under the GNU General Public License v3.0 at https://github.com/ChronoRoot/ChronoRoot2. This repository contains the full software implementation and comprehensive documentation to set up the system and utilizing the software.
  • The annotated image dataset used for training and validation contains 911 infrared images of Arabidopsis thaliana seedlings and 480 images of tomato with expert annotations for multiclass segmentation. This dataset is publicly available without restrictions at https://huggingface.co/datasets/ngaggion/ChronoRoot2. The dataset includes both raw images and their corresponding multi-class segmentation masks in .nii.gz format, as directly generated in the manual annotations made by our biologists. Scripts to convert to nnUNet’s standardized structure for 2D images are also provided within the GitHub repo, generating the correct splits to avoid mixing videos in training, validation and test partitions.
  • To facilitate reproducibility and allow users to familiarize themselves with the different analysis modules, we provide four complete demo datasets covering the scenarios presented in this paper. These datasets are available within the apptaniner and Docker images, via the ChronoRoot website, and as Supplementary Material to this manuscript:
    1. 1.
      Detailed Root Analysis: A video for RSA characterization of individual Arabidopsis plants.
    2. 2.
      Germination Screening: A video containing hundreds of seeds to test the germination analysis module.
    3. 3.
      Etiolation Screening: A video of seedlings grown in darkness for testing hypocotyl growth rates.
    4. 4.
      Tomato Comparison: A pair of videos of tomato illustrating cross-species capability.
6 Declarations
6.1 List of abbreviations
FPCA: Functional principal component analysis;
GPU: graphical processing unit;
IR: infrared;
LR: lateral root;
MR: main root;
PC: principal component;
ROI: region of interest;
RSA: root system architecture;
RSML: Root System Markup Language;
SORT: Simple online realtime tracking algorithm;
TMGR: Time of maximum germination rate;
TR: total root;
6.2 Ethical Approval
Not applicable.
6.3 Consent for publication
Not applicable.
6.4 Competing Interests
The authors declare that they have no competing interests.
6.5 Funding
6.6 Author’s Contributions
@@
  • F. Ariel, A. Barrios, F. Catulo, T. Blein and N. Boccardo conducted the biological experiments analyzed in the study
  • F. S. Rodriguez, F. E. Aballay, F. B. Catulo, M. F. Mammarella, M. F. Legascue, and S. N. Villarreal performed manual image annotation for training the Arabidopsis thaliana nnUNet model.
  • N. Boccardo and L. Santoro performed manual image annotation for training the tomato nnUNet model.
  • L. I. Pereyra-Bistrain and M. Benhamed identified the gene involved in root development, generated the tomato mutant and provided the plant material for Use Case 4.
@@
7 Acknowledgements
Not applicable.
8 Supplementary Materials
@@
Table S1 Detailed report of segmentation performance across plant organs for Arabidopsis thaliana.
Table S2 Detailed report of segmentation performance for tomato
@@
Table S1: Detailed report of segmentation performance across plant organs for arabidopsis thaliana.
Configuration Metric Complete Root Main Root Lateral Root Seed Hypocotyl Leaf Petiole
Residual DC 0.800±0.123 0.803±0.122 0.735±0.176 0.686±0.277 0.699±0.230 0.847±0.162 0.758±0.161
HD 6.658±9.703 6.649±9.895 8.948±11.217 14.811±16.148 10.323±13.220 5.169±7.701 7.160±12.339
Cp 0.930 0.933 0.895 - - - -
Cr 0.937 0.939 0.897 - - - -
Residual (Fast) DC 0.797±0.126 0.800±0.126 0.726±0.183 0.699±0.256 0.695±0.228 0.844±0.163 0.751±0.162
HD 7.162±10.081 7.243±10.218 9.931±10.178 14.850±15.882 10.051±13.968 5.348±9.297 7.592±13.285
Cp 0.926 0.929 0.901 - - - -
Cr 0.935 0.937 0.892 - - - -
Standard DC 0.798±0.119 0.800±0.119 0.720±0.195 0.696±0.281 0.695±0.240 0.844±0.165 0.758±0.156
HD 7.765±12.436 7.867±12.513 7.750±9.282 15.792±17.682 8.730±10.412 5.147±7.480 6.666±12.053
Cp 0.934 0.938 0.890 - - - -
Cr 0.937 0.938 0.915 - - - -
Standard (Fast) DC 0.793±0.123 0.796±0.123 0.711±0.203 0.698±0.276 0.687±0.245 0.840±0.166 0.753±0.155
HD 7.902±12.523 7.986±12.661 8.791±9.679 16.349±18.227 8.916±11.339 7.688±12.846 6.953±11.259
Cp 0.929 0.933 0.874 - - - -
Cr 0.936 0.936 0.901 - - - -
Table S2: Detailed report of segmentation performance across plant organs for tomato.
Training Configuration Metric Complete Root Main Root Lateral Root Seed Hypocotyl Aerial
Tomato Standard DC 0.836±0.178 0.840±0.178 0.643±0.294 0.866±0.202 0.844±0.197 0.778±0.225
HD 26.470±19.212 26.726±19.700 27.811±19.615 17.744±23.242 7.578±13.767 12.602±16.797
Cp 0.920±0.133 - - - - -
Cr 0.779±0.230 - - - - -
Tomato Standard (Fast) DC 0.805±0.194 0.812±0.188 0.615±0.289 0.884±0.163 0.834±0.201 0.768±0.235
HD 32.563±19.591 32.483±20.232 33.433±19.507 17.432±22.671 12.784±21.839 19.241±24.454
Cp 0.910±0.157 - - - - -
Cr 0.733±0.253 - - - - -
Tomato Residual DC 0.882±0.154 0.884±0.151 0.651±0.279 0.857±0.191 0.894±0.138 0.807±0.222
HD 19.676±20.152 15.238±16.119 23.138±20.741 18.005±21.034 6.021±11.986 11.156±17.283
Cp 0.908±0.153 - - - - -
Cr 0.868±0.189 - - - - -
Tomato Residual (Fast) DC 0.871±0.165 0.876±0.159 0.632±0.277 0.839±0.202 0.885±0.147 0.785±0.216
HD 23.438±20.654 18.147±18.024 24.354±20.547 18.172±20.642 6.290±11.842 13.965±18.417
Cp 0.904±0.159 - - - - -
Cr 0.853±0.197 - - - - -
Both Standard DC 0.853±0.173 0.862±0.162 0.639±0.324 0.885±0.159 0.850±0.186 0.791±0.218
HD 25.934±20.458 27.324±20.089 23.581±18.204 15.667±21.435 10.579±17.906 15.651±15.909
Cp 0.914±0.141 - - - - -
Cr 0.816±0.223 - - - - -
Both Standard (Fast) DC 0.842±0.181 0.849±0.172 0.638±0.312 0.887±0.150 0.851±0.187 0.771±0.234
HD 25.137±20.132 26.765±19.903 26.755±18.153 14.647±20.058 8.937±14.857 17.830±17.904
Cp 0.908±0.156 - - - - -
Cr 0.791±0.232 - - - - -
Both Residual DC 0.893±0.162 0.889±0.157 0.673±0.301 0.886±0.166 0.904±0.144 0.854±0.229
HD 9.943±13.860 11.355±14.100 15.416±16.647 11.604±17.573 5.976±12.757 14.552±23.488
Cp 0.916±0.138 - - - - -
Cr 0.896±0.177 - - - - -
Both Residual (Fast) DC 0.889±0.168 0.883±0.169 0.664±0.288 0.881±0.153 0.895±0.143 0.859±0.196
HD 10.218±13.685 12.090±13.817 17.317±18.033 12.134±17.901 6.445±13.448 13.454±19.100
Cp 0.905±0.165 - - - - -
Cr 0.899±0.172 - - - - -
Figure S4: Use Case 3 - FPCA Analysis.To complement the growth dynamics presented in Figure 5-B, we performed FPCA to summarize and compare the developmental trajectories of hypocotyl length, total root length, and projected area across genotypes G1, G2, and G3. The first two principal components accounted for over 90% of the total variance in each trait, capturing the main temporal patterns of growth. For each metric, the top panels show the mean trajectories with standard error bars for each genotype and a scatter plot of PC1 vs PC2. The middle and bottom rows illustrate the distribution of FPCA scores by group for PC1 and PC2, along with visual interpretations of each component. p-values: * < 0.05. ** < 0.001